INFRARED SPECTRA AND INTENSITIES OF WATER-AMMONIA COMPLEXES IN HELIUM DROPLETS

SUSUMU KUMA, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan; MIKHAIL N. SLIPCHENKO^{a}, Department of Chemistry, University of Southern California, Los Angeles, CA 90080, USA; \underline{TAKAMASA MOMOSE}, Department of Chemistry, University of British Columbia, Vancouver, BC V6T 1Z1, Canada; and ANDREY F. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90080, USA.

The infrared spectra of the H_{2}O-NH_{3} complexes in He droplets were measured in the range of the hydrogen stretching vibrational bands of water molecules. We observed two bands which were ascribed to the bonded and free OH stretching vibrations of water molecules. The infrared intensity ratio of the H_{2}O-NH_{3} complex and the single water molecule in this range was determined on the basis of the Poisson distribution of those abundances in He droplets. Infrared intensity of OH stretching vibrations in complexes was found to be 3 times larger that that of free molecules. The enhancement is ascribed to the formation of a hydrogen bond. Both the frequencies and intensities are compared to previous matrix isolation studies and theoretical calculations.

apresent address:1605 Gilman Hall, Iowa State University, Ames, IA 50011-3111, USA